# Curve fitting on a quantum annealer for an advanced navigation method

### v2

@@ Abstract
We explore the applicability of quantum annealing to the approximation task of curve fitting. To this end, we consider a function that shall approximate a given set of data points and is written as a finite linear combination of standardized functions, e.g., orthogonal polynomials. Consequently, the decision variables subject to optimization are the coefficients of that expansion. Although this task can be accomplished classically, it can also be formulated as a quadratic unconstrained binary optimization problem, which is suited to be solved with quantum annealing. Given the size of the problem stays below a certain threshold, we find that quantum annealing yields comparable results to the classical solution. Regarding a real-world use case, we discuss the problem to find an optimized speed profile for a vessel using the framework of dynamic programming and outline how the aforementioned approximation task can be put into play.
@@ I Introduction
The remainder of this article is organized as follows. In Section II, we formulate curve fitting as a QUBO and present our numerical results in Section III. After that, Section IV focuses on finding an optimized speed profile using dynamic programming, and we finally conclude in Section V.
@@ II Curve fitting as a QUBO
Let $X,Y\subset\mathbb{R}$, $n\in\mathbb{N}$, and $\{(x_{i},y_{i})\}_{i=0}^{n-1}\subset X\times Y$ the observed data points. The task is to find an optimal fit function $f^{\ast}\colon X\to Y$ in the sense that the sum of squares of the differences between the observed data $y_{i}$ and approximated values $f^{\ast}(x_{i})$ is minimized (method of least squares) [47]. With $\mathcal{F}$ denoting a suitable function space, e.g., =$\mathcal{F}=\{f\colon\mathbb{R}\to\mathbb{R}\mid f\text{\leavevmode\nobreak\ continuous}\}$, we are concerned with the optimization problem
@@
    $\Lambda_{j}(x)=\begin{dcases}\frac{x-\tilde{x}_{j-1}}{\tilde{x}_{j}-\tilde{x}_{j-1}}&\text{if\leavevmode\nobreak\ }x\in[\tilde{x}_{j-1},\tilde{x}_{j})\\[2.5pt] \frac{\tilde{x}_{j+1}-x}{\tilde{x}_{j+1}-\tilde{x}_{j}}&\text{if\leavevmode\nobreak\ }x\in[\tilde{x}_{j},\tilde{x}_{j+1})\\[2.5pt] 0&\text{otherwise}\end{dcases}$ (4)
@@ II.1 Classical solution
We aim to determine the expansion coefficients $c_{j}$ in Eq. 2 according to the least-squares principle as given by the optimization problem 1. With the abbreviations (where $j,k=0,\dotsc,m-1$)
@@
vectors $c=[c_{0},\dotsc,c_{m-1}]^{\top}\in\mathbb{R}^{m}$, $b=[b_{0},\dotsc,b_{m-1}]^{\top}\in\mathbb{R}^{m}$, and the symmetric matrix $W\in\mathbb{R}^{m\times m}$ with entries $W_{jk}$, the optimization problem upon inserting Eq. 2 into 1 now takes the form
@@ II.2 Quantum annealing-oriented approach
$\sigma_{r}=\begin{cases}-1&\text{if\leavevmode\nobreak\ }r=d-1\\[2.5pt] 1&\text{otherwise}\end{cases}$ (12)
accounts for two’s complement, i.e., the most significant bit enters the sum with a negative sign. For example, for $p=0$, Eq. 11 allows to represent all integers between $-(2^{d-1})$ and $12^{d-1}-$ 1. The quantities $\psi_{jr}$ are then the new decision variables represented by the qubit states of a quantum annealing device.
@@
Now, inserting Eq. 11 into 10 yields a new objective function depending on $\psi=[\psi_{00},\psi_{01},\dotsc]^{\top}\in\{0,1\}^{md}$, i.e., on all coefficients of the binary representation of the expansion coefficients; to wit
@@
where $\psi_{\mu}=\psi_{\hat{\mu}\bar{\mu}}$, $b_{\mu}^{\prime}=2^{\bar{\mu}-p+1}\sigma_{\bar{\mu}}b_{\hat{\mu}}$, and $W_{\mu\nu}^{\prime}=2^{\bar{\mu}+\bar{\nu}-2p}\sigma_{\bar{\mu}}\sigma_{\bar{\nu}}W_{\hat{\mu}\hat{\nu}}$ with $\hat{\mu}=\lfloor\mu/d\rfloor$ and $\bar{\mu}=\mu\bmod d$ (analogously for $\nu$). $b_{\mu}^{\prime}$ and $W_{\mu\nu}^{\prime}$ are the entries of the quantities $b^{\prime}\in\mathbb{R}^{md}$ and $W^{\prime}\in\mathbb{R}^{(md)\times(md)}$, respectively. These definitions allow us to write the objective function in a rather compact form:
@@
$Q=W^{\prime}-\operatorname{diag}(b_{0}^{\prime},\dotsc,b_{md-1}^{\prime})\in\mathbb{R}^{(md)\times(md)}\,.$ (17)
In a last step, the solution $\psi^{\ast}$ of the QUBO problem is plugged into Eq. 11 to obtain the actual, real-valued optimal coefficients $c_{j}^{\ast}=\sum_{r}\psi_{jr}^{\ast}\sigma_{r}2^{r-p}$.
Apparently, the number of required qubits, i.e., the size of the QUBO problem, depends only on the number of standardized functions $m$ used for the expansion of the fit function, see Eq. 2, and the number of digits $d$ used for the binary representation of the expansion coefficients. In other words, the size of the QUBO matrix $Q$ is independent of the number of the to-be-approximated data points $n$ (see Eq. 1); the latter enters only implicitly in the course of the determination of the entries of $Q$ (see Eqs. 5 and 6).
@@ III Numerical experiments
$Z_{\textup{Ising}}(s)\bigr{\rvert}_{s_{i}=2z_{i}-1}=Z_{\textup{QUBO}}(z)+\textup{const.}\,,$. , (19)
@@ III.1 Sample data
Before we continue with our actual results, we define the sample data $\{(x_{i},y_{i})\}_{i=0}^{n-1}$ (see Section II) that shall be used for our numerical experiments. The abscissa values are evenly spaced over the unit interval according to
@@
[…]
which are all distorted by noise $\varepsilon_{i}$. At each ordinate $y_{i}$, the latter is a random variable drawn from a normal distribution centered about zero with variance $(0.03)^{2}$, i.e., $\varepsilon_{i}\sim\mathcal{N}(\mu=0,\sigma^{2}=(0.03)^{2})$ for $i=0,\dotsc,n-1$.
@@
$\operatorname{ape}(a,b)=\frac{\lvert a-b\rvert}{\max\{\lvert a\rvert,\delta\}}\times 100$ (25)
with $a,b\in\mathbb{R}$, $\delta>0$ or the root mean squared error (RMSE)
@@ III.2 Triangular functions
We begin with triangular functions as defined in Eq. 4, $\phi_{j}=\Lambda_{j}$, i.e., the fit function is given by
$f(x)=c_{0}\Lambda_{0}(x)+\dotsb+c_{m-1}\Lambda_{m-1}(x)\,,$ (28)
and the least squares optimization problem thus explicitly reads
$\min_{c\in\mathbb{R}^{m}}\sum_{i=0}^{n-1}\,\Biggl{(}y_{i}-\sum_{j=0}^{m-1}c_{j}\Lambda_{j}(x_{i})\Biggr{)}^{2}\,.$ (29)
First, we consider the simple yet important case of linear regression. To this end, we choose $n=64$ and the linear sample data 21. Furthermore, we set $m=2$, which implies that the two supporting points of the triangular functions (see Eq. 4) are simply the first and last abscissa values 20 of the sample data, respectively, i.e., $\tilde{x}_{0}=x_{0}=0$ and $\tilde{x}_{1}=x_{n-1}=1$. This results in a fit function $f(x)=c_{0}\Lambda_{0}(x)+c_{1}\Lambda_{1}(x)$ that is a straight line across the whole abscissa interval $[0,1]$ of our sample data. For the corresponding QUBO formulation of the problem, we choose $d=10$ and $p=d-2=8$, which results in a QUBO matrix of size $md\times md=20\times 20$.
@@
Table 2: Same as Tab. 3 but with the linear sample data now given by $y_{i}=-x_{i}/4+1/3+\varepsilon_{i}$.
@@
Table 3: APE between the RMSE of the optimal classical solution $f_{\textup{cl}}^{\ast}$ and the optimal QUBO solutions $f_{\textup{tb}}^{\ast}$ and $f_{\textup{qa}}^{\ast}$ obtained via tabu search and quantum annealing, respectively, for different values of $n$. The underlying sample data set is the linear one (Eq. 21) and a linear regression is used for the approximation.
@@
In Tab. 3, we show our results for the optimal linear regression parameters $c_{0}^{\ast}$ and $c_{1}^{\ast}$ obtained classically by (naive) matrix inversion according to Eq. 9 and by solving the corresponding QUBO as given in Eq. 16; we also show the APE (see Eq. 25) of the optimal coefficients obtain via a QUBO with respect to the classical results. The QUBO is solved by means of a tabu search as well as by quantum annealing on a D-Wave quantum processing unit. [^2] For the tabu search, we used the MST2 multi-start tabu algorithm of Ref. [57]. Regarding quantum annealing for solving the QUBO, we used D-Wave’s Advantage2 system. In particular, we performed 100 runs with 100 reads each with an annealing time of 200 microseconds to collect potential solution candidates, yielding 10000 samples in total, and finally took the lowest-energy solution. This setup is used throughout the whole section whenever a QUBO problem is solved via tabu search or quantum annealing. Comparing the three solutions, we find that all agree up to the second decimal point. The QUBO solution found by the tabu search is particularly close to the exact classical solution, agreeing up to the third decimal point. Though the quantum annealing solution differs more noticeable, it is still in good agreement with the two other results. Furthermore, we find qualitatively similar results if we replace the numerical coefficients in the linear sample data 21 by other values, e.g., randomly drawn from $[-1,1]$. As an example, Tab. 3 shows the results for a different set of linear sample data.
@@
Thus, in the case of linear regression, our results indicate that formulating curve fitting as a QUBO problem indeed works and yields a solution in satisfying agreement with the classical one. They also suggest that quantum annealing in its current state is in principle able (within limits) to deliver promising and practically usable results for this continuous problem. [^3] Because of the probabilistic (or stochastic) nature of quantum annealing and the fact that the quantum annealing device is a real (i.e., non-ideal) system—thus susceptible to thermal fluctuations and interactions with its environment—subsequent calculations with the same input may yield slightly different results. Based on our experiences and observations in the course of this work, the quantum annealing results presented here seem to be accompanied with an overall error of about five to ten percent.
@@
It is furthermore instructive to investigate the influence of certain parameters on the solutions. In Tab. 3, we show the APE between the RMSE of the optimal classical solution $f_{\textup{cl}}^{\ast}$ and the optimal QUBO solutions $f_{\textup{tb}}^{\ast}$ and $f_{\textup{qa}}^{\ast}$ obtained via tabu search and quantum annealing, respectively, for different values of the number of sample data points $n$ (again for the case of linear regression with sample data 21). We conclude that the number of to-be-approximated data points has a negligible influence on the solution quality. This is expected because $n$ enters only implicitly in the course of determining the entries of the QUBO matrix (see Eqs. 5 and 6) while its size is given by $md\times md$, see Eq. 17, i.e., only determined by $m$ and $d$. We find the same result if we use our other sample data sets instead of the linear one. Therefore, we fix $n=64$ for the remainder of this section.
@@
In Fig. 1, we show the RMSE of the tabu (red squares) and quantum annealing (green diamonds) solutions $f_{\textup{tb}}^{\ast}$ and $f_{\textup{qa}}^{\ast}$, respectively, as a function of $d$, the number of digits that is used for the binary fixed-point format of the coefficients $c_{j}$, see Eq. 11, with $p=d-1$ fixed. Furthermore, we show the RMSE of the classical solution $f_{\textup{cl}}^{\ast}$, which is exact in the sense of numerical precision and thus independent of $d$. Here, we use the cubic sample data 23 and $m=4$, i.e., we are beyond linear regression and have a truly piecewise-linear fit of the form $f(x)=c_{0}\Lambda_{0}(x)+\dotsb+c_{3}\Lambda_{3}(x)$ that is not just a straight regression line. For $d\geq 8$, the tabu solutions are very close to the classical one and a further increase of $d$ has only a negligible impact. The same holds for the quantum annealing solutions, though they are always less accurate than the corresponding tabu results. Both are only identical provided the QUBO matrix is small enough, namely for $d\leq 5$, which corresponds to a QUBO matrix of size $20\times 20$ or smaller. In our opinion, however, the quantum annealing solutions for larger QUBO sizes are still in acceptable agreement with the tabu and classical ones. We find qualitatively similar results if we use the quadratic sample data or if the numerical coefficients of the cubic data are randomly drawn from $[-1,1]$. Therefore, $d=8$ (with $p=d-1$) seems to be sufficient for the binary fixed-point representation of the expansion coefficients $c_{j}$ regarding the cases considered in this work. Furthermore, $d\in\{6,7\}$ are reasonable compromises between accuracy and keeping the QUBO matrix as small as possible.
@@
Figure 1: RMSE of the $d$ tabu search (red squares, $f_{\textup{tb}}^{\ast}$) and quantum annealing (green diamonds, $f_{\textup{qa}}^{\ast}$) solutions as a function of $d$. The RMSE of the exact, $d$-independent classical solution $f_{\textup{cl}}^{\ast}$ is shown as a gray dashed line. Lines between data points are shown to guide the eye. The underlying sample data set is the cubic one (Eq. 23), and triangular functions with $m=4$ are used for the approximation.
@@
Last, we consider the case of varying $m$, i.e., how the solution quality depends on the number of triangular functions used for the curve fitting. This is an particularly interesting case because one usually chooses a sufficient numerical accuracy and simply increases $m$ to the point where the data is sufficiently good approximated. Curve fitting with triangular functions is well-suited for that because the result is a piecewise-linear approximation (a polygonal chain), which is often used to approximate more complex curves or if the functional form of the to-be-approximated data is not immediately obvious. In Fig. 2, we show the RMSE of the tabu (red squares), quantum annealing (green diamonds), and classical (gray circles) solutions as a function of $m$, where the sample data sets are the trigonometric (left diagram) and quadratic ones (right diagram); again, $n=64$, $d=8$, and $p=d-1=7$. First, we consider the left diagram. It shows the overall trend that the RMSE of the classical solution becomes smaller as $m$ increases. Initially, the tabu and quantum annealing solutions follow that trend. For $m\geq 10$ (which corresponds to a QUBO matrix of size $80\times 80$), however, the quantum annealing solution becomes worse with increasing $m$. The right diagram, though with a rather different underlying set of sample data (quadratic instead of trigonometric), shows the same qualitative behavior. Quantitatively, noticeable differences and a continuous worsening of the quantum annealing solutions with increasing $m$ already starts at $m=8$, which corresponds to a QUBO matrix size of $64\times 64$.
@@
We also tested other polynomial sample data as well as varied the numerical coefficients of the sample data used for Fig. 2 and found qualitatively similar results. This indicates that a QUBO size of about $60\times 60$ to $80\times 80$ seems to be the range above which D-Wave’s current quantum annealing hardware fails to find a satisfying solution of the QUBO formulation of least squares curve fitting with triangular functions.
@@
Table 4: Optimal expansion coefficients of the individual Chebyshev polynomials of a quadratic fit, obtained classically via matrix inversion and by solving a QUBO via tabu search as well as quantum annealing. Also shown are the APE of the optimal coefficients obtained from the QUBO with respect to the classical ones.
Classical (matrix inversion) $0.544598$ $-0.031799$ $0.499257$
QUBO (tabu search) $0.523438$ $-0.039063$ $0.484375$ $3.89$ $22.84$ $2.98$
QUBO (quantum annealing) $0.445313$ $-0.101563$ $0.421875$ $18.23$ $219.39$ $15.50$
Figure 3: RMSE of the tabu search (red squares, $f_{\textup{tb}}^{\ast}$) and quantum annealing (green diamonds, $f_{\textup{qa}}^{\ast}$) solutions as a function of $d$. The RMSE of the exact, $d$-independent classical solution $f_{\textup{cl}}^{\ast}$ is shown as a gray dashed line. Lines between data points are shown to guide the eye. The underlying sample data set is the cubic one (Eq. 23), and Chebyshev polynomials with $m=4$ are used for the approximation.

### v3

@@ Abstract
We explore the applicability of quantum annealing to the approximation task of curve fitting. To this end, we consider a function that shall approximate a given set of data points and is written as a finite linear combination of standardized functions, e.g., orthogonal polynomials. Consequently, the decision variables subject to optimization are the coefficients of that expansion. Although this task can be accomplished classically, it can also be formulated as a quadratic unconstrained binary optimization problem, which is suited to be solved with quantum annealing. Given the size of the problem stays below a certain threshold, we find that quantum annealing yields comparable results to the classical solution. Regarding a real-world use case, we discuss the problem to find an optimized speed profile for a vessel using the framework of dynamic programming and outline how the aforementioned approximation task can be put into play. Similar to the curve fitting task, our findings indicate that quantum annealing is currently only feasible if the routing problem is modeled sufficiently small and sparse.
@@ I Introduction
The remainder of this article is organized as follows. In Section˜II, we formulate curve fitting as a QUBO and present our numerical results in Section˜III. After that, Section˜IV focuses on finding an optimized speed profile using dynamic programming, and we finally conclude in Section˜V.
@@ II Curve fitting as a QUBO
Let $X,Y\subset\mathbb{R}$, $n\in\mathbb{N}$, and $\{(x_{i},y_{i})\}_{i=0}^{n-1}\subset X\times Y$ the observed data points. The task is to find an optimal fit function $f^{\ast}\colon X\to Y$ in the sense that the sum of squares of the differences between the observed data $y_{i}$ and approximated values $f^{\ast}(x_{i})$ is minimized (method of least squares) [47]. With $\mathcal{F}$ denoting a suitable function space, e.g., $\mathcal{F}=\{f\colon\mathbb{R}\to\mathbb{R}\mid f\text{\penalty 10000\ continuous}\}$, we are concerned with the optimization problem
@@
    $\Lambda_{j}(x)=\begin{dcases}\frac{x-\tilde{x}_{j-1}}{\tilde{x}_{j}-\tilde{x}_{j-1}}&\text{if\penalty 10000\ }x\in[\tilde{x}_{j-1},\tilde{x}_{j})\\[2.5pt] \frac{\tilde{x}_{j+1}-x}{\tilde{x}_{j+1}-\tilde{x}_{j}}&\text{if\penalty 10000\ }x\in[\tilde{x}_{j},\tilde{x}_{j+1})\\[2.5pt] 0&\text{otherwise}\end{dcases}$ (4)
@@ II.1 Classical solution
We aim to determine the expansion coefficients $c_{j}$ in Eq.˜2 according to the least-squares principle as given by the optimization problem 1. With the abbreviations (where $j,k=0,\dotsc,m-1$)
@@
vectors $c=[c_{0},\dotsc,c_{m-1}]^{\top}\in\mathbb{R}^{m}$, $b=[b_{0},\dotsc,b_{m-1}]^{\top}\in\mathbb{R}^{m}$, and the symmetric matrix $W\in\mathbb{R}^{m\times m}$ with entries $W_{jk}$, the optimization problem upon inserting Eq.˜2 into 1 now takes the form
@@ II.2 Quantum annealing-oriented approach
$\sigma_{r}=\begin{dcases}-1&\text{if\penalty 10000\ }r=d-1\\[2.5pt] 1&\text{otherwise}\end{dcases}$ (12)
accounts for two’s complement, i.e., the most significant bit enters the sum with a negative sign. For example, for $p=0$, Eq.˜11 allows to represent all integers between $(2^{d-1})$ and $2^{d-1}-1$. The quantities $\psi_{jr}$ are then the new decision variables represented by the qubit states of a quantum annealing device.
@@
Now, inserting Eq.˜11 into 10 yields a new objective function depending on $\psi=[\psi_{00},\psi_{01},\dotsc]^{\top}\in\{0,1\}^{md}$, i.e., on all coefficients $\psi_{jr}$ of the binary representation of the expansion coefficients; to wit
@@
where we introduced $\psi_{\mu}=\psi_{\hat{\mu}\bar{\mu}}$, $b_{\mu}^{\prime}=2^{\bar{\mu}-p+1}\sigma_{\bar{\mu}}b_{\hat{\mu}}$, and $W_{\mu\nu}^{\prime}=2^{\bar{\mu}+\bar{\nu}-2p}\sigma_{\bar{\mu}}\sigma_{\bar{\nu}}W_{\hat{\mu}\hat{\nu}}$ with $\hat{\mu}=\lfloor\mu/d\rfloor$, $\bar{\mu}=\mu\bmod d$ (analogously for $\nu$). $b_{\mu}^{\prime}$ and $W_{\mu\nu}^{\prime}$ are the entries of the quantities $b^{\prime}\in\mathbb{R}^{md}$ and $W^{\prime}\in\mathbb{R}^{(md)\times(md)}$, respectively. These definitions allow us to write the objective function in a rather compact form:
@@
$Q=W^{\prime}-\operatorname{diag}(b_{0}^{\prime},\dotsc,b_{md-1}^{\prime})\,.$ (17)
In a last step, the solution $\psi^{\ast}$ of the QUBO problem is plugged into Eq.˜11 to obtain the actual, real-valued optimal coefficients $c_{j}^{\ast}=\sum_{r}\psi_{jr}^{\ast}\sigma_{r}2^{r-p}$.
Apparently, the number of required qubits, i.e., the size of the QUBO problem, depends only on the number of standardized functions $m$ used for the expansion of the fit function, see Eq.˜2, and the number of digits $d$ used for the binary representation of the expansion coefficients:
$Q\in\mathbb{R}^{(md)\times(md)}\,.$ (18)
In particular, the size of the QUBO matrix is independent of the number of the to-be-approximated data points $n$ (see Eq.˜1); the latter enters only implicitly in the course of calculating the entries of $Q$ (see Eqs. 5 and 6).
@@ III Numerical experiments
$Z_{\textup{Ising}}(s)\bigr\rvert_{s_{i}=2z_{i}-1}=Z_{\textup{QUBO}}(z)+\textup{const.}\,$, (20)
@@ III.1 Sample data
Before we continue with our actual results, we define the sample data $\{(x_{i},y_{i})\}_{i=0}^{n-1}$ (see Section˜II) that shall be used for our numerical experiments. The abscissa values are evenly spaced over the unit interval according to
@@
$\textup{linear:}\quad y_{i}$ $=\frac{1}{2}x_{i}+1+\varepsilon_{i}\,$, (22)
$\textup{quadratic:}\quad y_{i}$ $=\frac{3}{4}x_{i}^{2}+\varepsilon_{i}\,$, (23)
$\textup{cubic:}\quad y_{i}$ $=\frac{3}{4}x_{i}^{3}+\frac{1}{4}x_{i}+\varepsilon_{i}\,$, (24)
$\textup{trigonometric:}\quad y_{i}$ $=\sin(2\pi x_{i})\cos(2\pi x_{i})+\varepsilon_{i}\,$, (25)
which are all distorted by artificial noise $\varepsilon_{i}$. At each ordinate $y_{i}$, the latter is a random variable drawn from a normal distribution centered about zero with variance $(0.03)^{2}$, i.e., $\varepsilon_{i}\sim\mathcal{N}(\mu=0,\sigma^{2}=(0.03)^{2})$.
@@
$\operatorname{ape}(a,b)=\frac{\lvert a-b\rvert}{\max\{\lvert a\rvert,\delta\}}\times 100\,$, (26)
where $a,b\in\mathbb{R}$ and $\delta>0$, or the root mean squared error (RMSE)
@@ III.2 Triangular functions
Table 2: Same as Tab.˜1 but with the linear sample data now given by $y_{i}=-x_{i}/4+1/3+\varepsilon_{i}$.
@@
Table 3: APE between the RMSE of the optimal classical solution $f_{\textup{cl}}^{\ast}$ and the optimal QUBO solutions $f_{\textup{tb}}^{\ast}$ and $f_{\textup{qa}}^{\ast}$ obtained via tabu search and quantum annealing, respectively, for different values of $n$. The underlying sample data set is the linear one (Eq.˜22) and a linear regression is used for the approximation.
@@
We begin with triangular functions as defined in Eq.˜4, $\phi_{j}=\Lambda_{j}$, i.e., the fit function is given by
$f(x)=c_{0}\Lambda_{0}(x)+\dotsb+c_{m-1}\Lambda_{m-1}(x)\,$, (29)
and the least squares optimization problem thus explicitly reads
$\min_{c\in\mathbb{R}^{m}}\sum_{i=0}^{n-1}\left(y_{i}-\sum_{j=0}^{m-1}c_{j}\Lambda_{j}(x_{i})\right)^{2}\,.$ (30)
First, we consider the simple yet important case of linear regression. To this end, we choose $n=64$ and the linear sample data 22. Furthermore, we set $m=2$, which implies that the two supporting points of the triangular functions (see Eq.˜4) are simply the first and last abscissa values 21 of the sample data, respectively, i.e., $\tilde{x}_{0}=x_{0}=0$ and $\tilde{x}_{1}=x_{n-1}=1$. This results in a fit function $f(x)=c_{0}\Lambda_{0}(x)+c_{1}\Lambda_{1}(x)$ that is a straight line across the whole abscissa interval $[0,1]$ of our sample data. For the corresponding QUBO formulation of the problem, we choose $d=10$ and p=d2=8p=d-2=8, which results in a QUBO matrix of size $md\times md=20\times 20$.
@@
In Tab.˜1, we show our results for the optimal linear regression parameters $c_{0}^{\ast}$ and $c_{1}^{\ast}$ obtained classically by (naive) matrix inversion according to Eq.˜9 and by solving the corresponding QUBO as given in Eq.˜16; we also show the APE (see Eq.˜26) of the optimal coefficients obtain via a QUBO with respect to the classical results. The QUBO is solved by means of a tabu search as well as by quantum annealing on a D-Wave quantum processing unit (QPU). [^2] For the tabu search, we used the MST2 multi-start tabu algorithm of Ref. [57]. Regarding quantum annealing for solving the QUBO, we used D-Wave’s Advantage2 system. In particular, we performed 100 runs with 100 reads each with an annealing time of 200 microseconds to collect potential solution candidates, yielding 10000 samples in total, and finally took the lowest-energy solution. This setup is used throughout the whole section whenever a QUBO problem is solved via tabu search or quantum annealing. Comparing the three solutions, we find that all agree up to the second decimal point. The QUBO solution found by the tabu search is particularly close to the exact classical solution, agreeing up to the third decimal point. Though the quantum annealing solution differs more noticeable, it is still in good agreement with the two other results. Furthermore, we find qualitatively similar results if we replace the numerical coefficients in the linear sample data 22 by other values, e.g., randomly drawn from [1,1][-1,1]. As an example, Tab.˜3 shows the results for a different set of linear sample data.
@@
Thus, in the case of linear regression, our results indicate that formulating curve fitting as a QUBO problem indeed works and yields a solution in satisfying agreement with the classical one. They also suggest that quantum annealing in its current state is in principle able (within limits) to deliver promising and practically usable results for this continuous problem. [^3] Because of the probabilistic (or stochastic) nature of quantum annealing and the fact that the QPU is a real (i.e., non-ideal) system—thus susceptible to thermal fluctuations and interactions with its environment—subsequent calculations with the same input may yield slightly different results. Based on our experiences and observations in the course of this work, the quantum annealing results presented here seem to be accompanied with an overall error of about five to ten percent.
@@
It is furthermore instructive to investigate the influence of certain parameters on the solutions. In Tab.˜3, we show the APE between the RMSE of the optimal classical solution $f_{\textup{cl}}^{\ast}$ and the optimal QUBO solutions $f_{\textup{tb}}^{\ast}$ and $f_{\textup{qa}}^{\ast}$ obtained via tabu search and quantum annealing, respectively, for different values of the number of sample data points $n$ (again for the case of linear regression with sample data 22). We conclude that the number of to-be-approximated data points has a negligible influence on the solution quality. This is expected because $n$ enters only implicitly in the course of determining the entries of the QUBO matrix (see Eqs. 5 and 6) while its size is given by $md\times md$, see Eq.˜17, i.e., only determined by $m$ and $d$. We find the same result if we use our other sample data sets instead of the linear one. Therefore, we fix $n=64$ for the remainder of this section.
@@
In Fig.˜1, we show the RMSE of the tabu (red squares) and quantum annealing (green diamonds) solutions $f_{\textup{tb}}^{\ast}$ and $f_{\textup{qa}}^{\ast}$, respectively, as a function of $d$, the number of digits that is used for the binary fixed-point format of the coefficients $c_{j}$, see Eq.˜11, with p=d1p=d-1 fixed. Furthermore, we show the RMSE of the classical solution $f_{\textup{cl}}^{\ast}$, which is exact in the sense of numerical precision and thus independent of $d$. Here, we use the cubic sample data 24 and $m=4$, i.e., we are beyond linear regression and have a truly piecewise-linear fit of the form $f(x)=c_{0}\Lambda_{0}(x)+\dotsb+c_{3}\Lambda_{3}(x)$ that is not just a straight regression line. For $d\geq 8$, the tabu solutions are very close to the classical one and a further increase of $d$ has only a negligible impact. The same holds for the quantum annealing solutions, though they are always less accurate than the corresponding tabu results. Both are only identical provided the QUBO matrix is small enough, namely for $d\leq 5$, which corresponds to a QUBO matrix of size $20\times 20$ or smaller. In our opinion, however, the quantum annealing solutions for larger QUBO sizes are still in acceptable agreement with the tabu and classical ones. We find qualitatively similar results if we use the quadratic sample data or if the numerical coefficients of the cubic data are randomly drawn from [1,1][-1,1]. Therefore, $d=8$ (with p=d1p=d-1) seems to be sufficient for the binary fixed-point representation of the expansion coefficients $c_{j}$ regarding the cases considered in this work. Furthermore, $d\in\{6,7\}$ are reasonable compromises between accuracy and keeping the QUBO matrix as small as possible.
@@
Figure 1: RMSE of the tabu search (red squares, $f_{\textup{tb}}^{\ast}$) and quantum annealing (green diamonds, $f_{\textup{qa}}^{\ast}$) solutions as a function of $d$. The RMSE of the exact, $d$-independent classical solution $f_{\textup{cl}}^{\ast}$ is shown as a gray dashed line. Lines between data points are shown to guide the eye. The underlying sample data set is the cubic one (Eq.˜24), and triangular functions with $m=4$ are used for the approximation.
@@
Last, we consider the case of varying $m$, i.e., how the solution quality depends on the number of triangular functions used for the curve fitting. This is an particularly interesting case because one usually chooses a sufficient numerical accuracy and simply increases $m$ to the point where the data is sufficiently good approximated. Curve fitting with triangular functions is well-suited for that because the result is a piecewise-linear approximation (a polygonal chain), which is often used to approximate more complex curves or if the functional form of the to-be-approximated data is not immediately obvious. In Fig.˜2, we show the RMSE of the tabu (red squares), quantum annealing (green diamonds), and classical (gray circles) solutions as a function of $m$, where the sample data sets are the trigonometric (left diagram) and quadratic ones (right diagram); again, $n=64$, $d=8$, and p=d1=7p=d-1=7. First, we consider the left diagram. It shows the overall trend that the RMSE of the classical solution becomes smaller as $m$ increases. Initially, the tabu and quantum annealing solutions follow that trend. For $m\geq 10$ (which corresponds to a QUBO matrix of size $80\times 80$), however, the quantum annealing solution becomes worse with increasing $m$. The right diagram, though with a rather different underlying set of sample data (quadratic instead of trigonometric), shows the same qualitative behavior. Quantitatively, noticeable differences and a continuous worsening of the quantum annealing solutions with increasing $m$ already starts at $m=8$, which corresponds to a QUBO matrix size of $64\times 64$.
@@
We also tested other polynomial sample data as well as varied the numerical coefficients of the sample data used for Fig.˜2 and found qualitatively similar results. This indicates that a QUBO size of about $60\times 60$ to $80\times 80$ seems to be the range above which D-Wave’s current quantum annealing hardware fails to find a satisfying solution of the QUBO formulation of least squares curve fitting with triangular functions.